OpenStudy (anonymous):

Part 1 Is the following equation true? Part 2 Use complete sentences to explain the properties used in making your decision. 4 log 3^ (x + 1/2) log 3^ (y - 2) log 3 z = log 3 (x^4sqrt (y)/z^2

OpenStudy (ja1):

Oh my... not for me sorry I don't d this level math yet :)

OpenStudy (debbieg):

Is this the equation??\[4\log_{3}(x+\frac{1}{2}) \log_{3}(y-2) \log_{3}z=\frac{\log_{3}x^4\sqrt{y}}{z^2} \]

OpenStudy (debbieg):

or this?\[4\log_{3}x+\frac{1}{2} \log_{3}y-2\log_{3}z=\frac{\log_{3}x^4\sqrt{y}}{z^2} \]